OpenStudy (anonymous):

(6a + b)^2 =36a^2+ b^2?

OpenStudy (uri):

Correct.

OpenStudy (anonymous):

how do you know its been three secs???

OpenStudy (uri):

6*6=36 a+a =a^2 b+b=b^2 I need 2 sec to look at it :p

OpenStudy (anonymous):

ok sorry your pretty quick

OpenStudy (uri):

It's okay.. :)